10 gram of MNO2 on reaction with HCl forms 2.24 litre of Cl2 at NTP ,the percentage impurity of MNO2 is

Grade:12

1 Answers

ANIKET RAJ
5 years ago
first write and balance the equation
MnO+4HCL---MnCl2+2H2O
A/Q
Moles of Cl2=2.24/22.4==0.1moles
now,
A/equation 
1mole cl2 react with 1 mole mno2
mole of mno2=0.1
wt. of mno2=0.1*55=5.5g
 
impurity=10g-5.5g=4.5g
%=4.5/10 *100=45%ans